# General relativity and the double-zero eigenvalue

### v1

@@ 5.1 Friedmann universes
In this subsection we follow closely [19] and write the standard Friedmann equations as a 2-dimensional system of evolution equations for the Hubble parameter $H$ and the normalized density $\Omega$. Then in the Figure 1 (left) below we sketch the solution set, i.e., the $(\Omega,\mu)$-plane, where $\mu=3\gamma-2$, $\gamma$ being the fluid parameter of a fluid with equation of state $p=(\gamma-1)\rho$.
Figure 1: The standard (left) and bifurcation (right) density parameter evolution in Friedmann cosmology compared.
This is the picture one knows from the case of standard Friedmann cosmology.
@@
We then apply the bifurcation sequence steps of Section 4 and arrive at the versal unfolding $Z^{\prime}=\bar{\mu}-Z^{2}$, and the bifurcation diagram in Figure 1 (right) for the versally unfolded density $Z=1/2-\Omega$, where $\bar{\mu}=1/4-\sigma/\mu$, with $\mu\neq 0$, and $\sigma$ is the unfolding parameter which describes physically the closeness of the models to the unperturbed Friedmann universe (this can be made precise).
We note that in this problem both the parameter space and the phase space are one-dimensional, and the bifurcation takes place at the origin of the $(Z,\mu)$-plane.
@@
This leads to physical effects associated with the appearance of the new branch of fixed points, namely, the parabola shown in the Fig. 1 (right). Several features of the standard cosmological picture are now replaced by novel properties stemming from this Figure.
@@
For example, the equilibrium solutions 0 and 1 of the evolution equation for the density parameter $\Omega$ in standard cosmology are now absent and are replaced by new perturbed equilibria, i.e., pairs of the form ±$\pm\sqrt{\bar{\mu}}$. This has various implications that will be considered elsewhere.
We also note in passing the following important application of the versally unfolded Friedmann evolution. Although the original Friedmann equations cannot lead to domain synchronization in the dynamical sense considered in [19], Sect. 8, the versal unfolding $Z^{\prime}=\bar{\mu}-Z^{2}$ indeed synchronizes the universe asymptotically, either to the future or the past. This in turn leads to a novel approach to the horizon problem, as explained in detail in [19].
@@
Under parameter variation the novel solutions ((un)stable with the (minus) plus sign) appear on the parabola for positive values of the parameter (the vertical lines being the $Z$-phase spaces), coalesce at the origin and disappear for negative values of the parameter.
These novel forms of evolution are collectively described by what is known as a saddle-node bifurcation, the simplest kind of change of forms in bifurcation theory (the phase portraits for the saddle-node are very standard and not shown presently).
The perturbed equilibria possess further novel properties which we only briefly mention here (we direct the reader to [19] for more details and complete proofs):
  1. 1.
    unlike the situation in standard cosmology no singularities exist in this case,
  2. 2.
    the plus-sign equilibrium in each pair describes stable accelerating universes for a negative unfolding parameter (i.e., when $\bar{\mu}>1/4$ in Fig. 1, right), still all energy conditions are nevertheless satisfied (that is for $\gamma>2/3$).
@@ 5.2 The Oppenheimer-Snyder black hole
In this subsection we closely follow [20], Sect. 7 and parts of Sect. 8. The standard Oppenheimer-Snyder (‘OS’) result describing the optical disconnection of the exterior spacetime and the formation of a singularity at the centre of the black hole in a finite time follows from their solution $e^{x}=(F\tau+G)^{4/3}$, where $x$ stands here for the OS comoving function $\omega$, $F,G$ are arbitrary functions of the radial coordinate $R$, and one uses the Schwarzschild metric in comoving coordinates.
The Einstein equations for this problem are equivalent to the system $\ddot{x}+(3/4)\dot{x}^{2}=0$, for which the phase portrait orbits show the characteristic OS diverging behaviour (cf. [20]).
It is an instructive initial conclusion borne out of an application of the bifurcation/singularity theory steps described above, that:
@@
The versal unfolding for the OS problem thus requires the inclusion of 3rd-order terms for determinacy, and the problem is of codimension two.
We note the planar parameter spaces here rather than the parameter lines we met in the Friedmann problem in Fig. 1, and the two parameters, namely, deviation from spherical symmetry and rotation, make up the bifurcation diagrams for the deformations. These are shown in the two figures 2, 3 below corresponding to positive and negative modular coefficients respectively.
In these diagrams, we imagine a point moving in parameter space on a small circle around the origin and consider the corresponding motion of a phase point moving on some orbit in one of the phase spaces shown. As the parameter point moves around in parameter space, the phase portraits smoothly transfigure and the phase point drags away from one phase diagram to the next and gradually passes through the various metamorphoses of the different topological states (i.e., phase portraits in this case) as shown.
Figure 2: The bifurcation diagram for the Oppenheimer-Snyder-system, positive moduli case.
Figure 3: The bifurcation diagram for the Oppenheimer-Snyder-system, negative moduli case.
Three basic features of these diagrams are:
@@
    the formation of stable closed orbits for various values of the parameters in the left part of the diagram.
@@
    upon parameter variation, the singularities that would have possibly resulted from trapped surface formation and the OS solution seem to be stably bypassed. The system instead smoothly transfigures its phase portraits as shown, and any phase point is able to follow other orbits thus easily avoiding staying too long on any runaway orbit.
This is a very rich situation with novel features unlike anything in the original OS analysis. In addition, we note that there are additional global bifurcations not shown in these figures.
This constitutes a clear prediction of the physics associated with the versal unfolding in this problem. Under parameter variation, a trapped surface may initially form but such a configuration is always unstable and transfigures into something else precisely defined depending on the position of the parameter point in the stratified space (we recall a stratum as a region in parameter space between two consequent bifurcation sets, e.g., between say a bifurcation axis and a bifurcation line, etc).
@@ 5.3 Spacetime metamorphoses
In this problem we revisit the evolutionary aspects of the analysis that leads to the singularities in gravitational collapse to a black hole and in a cosmological setting, cf. [20], Sections 5, 6. In terms of the Komar-Landau-Raychaudhuri and the Newman-Penrose-Sachs equations, this analysis constitutes another manifestation of the power of bifurcation theory.
@@
Firstly, it allows for a complete analysis of the nonlinear feedback and coupling effects between the expansion, shear, and vorticity of a geodesic congruence in spacetime taking into full account all degeneracies of this problem. There are two cases, namely, the convergence-shear and the convergence-vorticity systems (subsystems of the Sachs optical equations), and we shall discuss presently the main ideas and results of the first case (this is in fact the only case considered in the classical references).
Secondly, one can now revisit the classical prediction about the existence of a focussing state. In the classical references this question is of course decided based on the analysis of the Raychaudhuri inequality $\rho^{\prime}\geq\rho^{2}$ for the convergence of the spacetime congruence leading to the prediction of spacetime singularities.
Applying the methodology of Sec. 4 to the convergence-shear and the convergence-vorticity dynamical equations leads to codimension-two problems and the final bifurcation diagrams are shown in Figs. 4, 5 respectively. Here the parameters $\mu_{1}$ and $\mu_{2}$ are related to the full projections of the Ricci curvature along a timelike vector and the Weyl curvature along a null tetrad (and rotation parameter respectively for the vorticity case), and the phase space variables are the convergence and the shear (resp. vorticity) of the geodesic congruence.
Figure 4: The complete bifurcation diagram for the NPR-system.
Figure 5: The complete bifurcation diagram for the convergence-vorticity system.
@@
As the parameter point moves on a small circle around the origin in the plane, the phase portraits smoothly transfigure to one another, and a phase point rides around from one phase portrait to the next. For instance, the transition from $\chi$ to $\varepsilon_{-}$ to $\pi$ to $\tau_{-}$ to $\beta$ strata (or for the opposite direction) is accompanied by ‘metamorphoses’ of non-equivalent phase diagrams as shown containing 0 to 1 to 2 to 3 equilibria respectively (or in reverse).
@@
In the emerging physics of the versal unfolding (instead of the standard physical conclusions based on the Raychaudhuri inequality) there are several novel physical effects, and we only note here two, namely, the possibility of spacetime metamorphoses and the ‘ghost effect’ associated with a square root scaling law. Both these effects make the existence of an all-encompassing singular state in general relativity less plausible.
Had the bifurcation diagram in Fig. 4 contained only the $\chi$-stratum (i.e., only the right half-space), there would be one phase portrait and nothing else as is the case with the Raychaudhuri inequality. In that case, there would be of course no possibility for a metamorphosis because the focussing state leading to a singularity would be an inevitable feature of the dynamics (corresponding to the diverging orbits in the $\chi$-phase portrait, the only kind of orbit there).
@@
However, here the situation is rather different: the system exhibits two kinds of bifurcations, namely, a pair of saddle-nodes dominated by the convergence $\rho$, and a pair of pitchfork bifurcations dominated by the shear $\sigma$ (recall that while the saddle-node bifurcation is the basic mechanism by which equilibria are created and destroyed, a pitchfork comes in two types, the supercritical by which an unstable equilibrium slowly decays into two new stable ones, and the subcritical where the system destabilizes).
A saddle-node bifurcation takes the system from the region of attractive gravity $\mu_{1}>0$ where all the energy conditions hold, to the repulsive gravity $\mu_{1}<0$, and back, while a pitchfork occurs in the repulsive gravity region of the parameter space (supercritical for $o\to\tau_{+}\to\beta$, subcritical for $\pi\to\tau_{-}\to\beta$ and back). The evolution of the system is characterized by the combined effects of all these bifurcations.
A second notable feature of the bifurcation dynamics, the ghost or ‘bottleneck’ effect, is related to the saddle-node evolution from and to the $\chi$-stratum, the ‘singularity-forming region’, and implies that the system delays its evolution upon entering the $\chi$-stratum with the delay time scaling as $\mu_{1}^{-1/2}$, a parameter-dependence of the time to the ‘singularity’ as a square-root scaling law, also absent in standard treatments.

### v2

@@ 3.2 Friedmann universes
In this subsection, we review the main results of the application of bifurcation theory to the Friedmann equations describing simple fluid-filled universes. We follow closely [21] and consider the standard Friedmann equations as a 2-dimensional system of evolution equations for the Hubble parameter $H$ and the normalized density $\Omega$, using a dimensionless time variable $\tau$ defined by $dt/d\tau=1/H$, namely,
$\frac{d\Omega}{d\tau}$ $=-\mu\Omega+\mu\Omega^{2}$ (3.1)
$\frac{dH}{d\tau}$ $=-H-\frac{1}{2}\mu\Omega H.$ (3.2)
This system has a number of non-hyperbolic equilibria of which we shall only consider the non-hyperbolic Milne state, namely, the $\Omega$-axis, EQ-1: $(\Omega,H,\mu)=(\Omega,0,0)$ (for the full analysis of the dispersive steady states we refer to [21], Sects. 5-7). It is important to realize that this system is not a true 2-dimensional system near each of the non-hyperbolic equilibria, but the dynamics is consequently restricted to a 1-dimensional invariant manifold along the dispersive dimension. Therefore a centre manifold reduction near EQ-1 gives the following dynamics on the centre manifold $W^{c}_{loc}$ (which in this case turns out to be the $\Omega$-axis) [21],
$\Omega^{\prime}$ $=-\mu\Omega+\mu\Omega^{2}$ (3.3)
$\mu^{\prime}$ $=0.$ (3.4)
@@
This analysis leads to Figure 1 (left) where we have sketched the solution set, i.e., the $(\Omega,\mu)$-plane, where $\mu=3\gamma-2$, $\gamma$ being the fluid parameter of a fluid with equation of state $p=(\gamma-1)\rho$.
Figure 1: Comparison of the density parameter evolution on the parametrized centre manifold (i.e., on $H=0$, $\mu$ being a new dependent variable) in Friedmann cosmology: standard case (left) vs. bifurcation case (right). On the left we see the phase-parameter diagram of the system (3.3)-(3.4), with the $\Omega$ phase spaces being the lines orthogonal to the $\mu$-axis. The red half-axes show instability and the horizontal black ones signify stability. On the right we see the cosmological saddle-node bifurcation diagram of the system (3.7) for the evolution of the universe. There are two new equilibria – the stable node and the saddle – the stable orbit is in the first quadrant. Phase lines are shown for two different $\bar{\mu}$-values.
We then apply the bifurcation sequence steps of Section 2 and arrive at the versal unfolding, [21],
$\frac{d\Omega}{dT}=\Omega(\Omega-1)+\nu,$ (3.5)
where,
$\nu=\frac{\sigma}{\mu},$ (3.6)
is the unfolding-fluid parameter and we have rescaled the time to $T=\mu\tau$. In terms of the versally unfolded density $Z=1/2-\Omega$, that is with $\Omega$ being a solution of (3.5), we arrive at the following form of the versal unfolding (3.5),
$\frac{dZ}{dT}=\bar{\mu}-Z^{2},\quad\bar{\mu}=\frac{1}{4}-\nu.$ (3.7)
@@
This is then the normal form of a saddle-node bifurcation with parameter $\bar{\mu}$, and bifurcation point at $(Z,\bar{\mu})=(0,0)$, and so includes the effects of perturbing terms of all other orders. The bifurcation diagram is shown in Figure 1 (right), where $\bar{\mu}=1/4-\sigma/\mu$, with $\mu\neq 0$, and $\sigma$ is the unfolding parameter which describes physically the closeness of the models to the unperturbed Friedmann universe (this can be made precise).
@@
We note that in this problem both the parameter space and the phase space are one-dimensional, and the bifurcation takes place at the origin of the $(Z,\bar{\mu})$-plane. This leads to physical effects associated with the appearance of the new branch of fixed points, namely, the parabola shown in the Fig. 1 (right). We note that these solutions have no counterpart in the standard picture (which only has the $\Omega=0,1$ equilibria), and so several features of the standard cosmological picture are now replaced by novel properties stemming from this Figure.
@@
For example, the equilibrium solutions 0 and 1 of the evolution equation for the density parameter $\Omega$ in standard cosmology are now absent and are replaced by new perturbed equilibria, i.e., pairs of the form $Z_{\pm}=\pm\sqrt{\bar{\mu}}$. This has various interesting implications.
3.2.1 Three applications of the versal FRW equations
@@
First, under parameter variation the novel solutions ((un-)stable with the (minus) plus sign) appear on the parabola for positive values of the parameter (the vertical lines being the $Z$-phase spaces), coalesce at the origin and disappear for negative values of the parameter (the phase portraits for the saddle-node and not shown presently). The perturbed equilibria possess novel properties (we direct the reader to [21] for more details and complete proofs), for instance, unlike the situation in standard cosmology no singularities exist in this case.
Another novel application of the versally unfolded Friedmann evolution that is impossible with the original Friedmann dynamics governed by Eq. (3.4) relates to the question of synchronization, cf. [21] (for other related references on the general problem of cosmological sync, we refer to [87], [88], [86]). This is the general question as to whether one or more causally disconnected Friedmann domains can sync during their evolution if they were not so synced initially (for a single domain $A$, sync means that if it is initially synced it remains so during its evolution). This can be made precise by introducing the sync function $\omega(\Omega,\mu)$ which measures differences in the density profiles of Friedmannian domains during evolution, and we say that A domain remains synced during evolution provided $\lim_{\tau\rightarrow\infty}\omega(\Omega|_{A},\mu)=0.$ It is a basic result that although the original Friedmann equations cannot lead to domain synchronization in this dynamical sense, the versal unfolding $Z^{\prime}=\bar{\mu}-Z^{2}$ indeed synchronizes the universe asymptotically, either to the future or the past. This in turn leads to a novel solution to the horizon problem, as explained in detail in [21].
Still another application of the versal family (3.7) is related to the existence of novel accelerating solutions absent in the standard model, cf. [25]. The construction of such solutions is an elaborated process and so we shall presently restrict our discussion to only the first step: What is the physical significance of the unfolding parameter $\sigma$ that appears in the versal unfolding equation (3.5)and in the $Z_{\pm}$ equilibria? The latter can be written in terms of the $\Omega$ function, that is,
$\Omega_{\pm}=\frac{1}{2}\left(1\mp\sqrt{1-4\nu}\right),\quad\nu<\frac{1}{4},$ (3.8)
and this implies that all solutions are stably attracted by $\Omega_{+}$ in the future, and by $\Omega_{-}$ in the past, and we find that the $\Omega_{+}$ equilibrium describes universes that are always open, while the $\Omega_{-}$ equilibria can be open (resp. closed) according to $\nu>0$ (resp. $\nu<0$). Writing the versal unfolding equation (3.5) in terms of the standard energy density $\rho$, we find that it leads to the form
$\dot{\rho}+3H(\rho+p)=3\sigma H^{3},$ (3.9)
which means that a nonzero $\sigma$ determines an out-of-equilibrium evolution and the non-constancy of entropy, with an entropic force term proportional to $\sigma H^{3}$. It turns out that this is key to understanding the nature of the versal family (3.7) and has important implications for the $\Omega_{\pm}$ steady states. We note only the following: the deceleration parameter $q=\frac{1}{2}\mu\Omega_{+},$ implies that there are accelerating solutions with either increasing or decreasing entropy, namely, $q<0$ even when $\mu>0$ and the strong energy condition is satisfied and any $\sigma>0$, or $<0$. Analogous results hold for the $\Omega_{-}$ states, cf. [25]. We shall refrain from presenting the complete arguments which require an extensive discussion of various aspects of the bifurcation properties of this model [25].
@@ 3.3 The Oppenheimer-Snyder black hole
In this subsection we closely follow [22], Sect. 7 and parts of Sect. 8. The standard Oppenheimer-Snyder (‘OS’) result describing the optical disconnection of the exterior spacetime and the formation of a singularity at the centre of the black hole in a finite time follows from their solution $e^{x}=(F\tau+G)^{4/3}$, where $x$ stands here for the OS comoving function $\omega$, $F,G$ are arbitrary functions of the radial coordinate $R$, and one uses the Schwarzschild metric in comoving coordinates. The Einstein equations for this problem are equivalent to the system [89],
$\ddot{x}+(3/4)\dot{x}^{2}=0,$ (3.10)
for which the phase portrait orbits show the characteristic OS diverging behaviour (cf. [22], Fig. 12). Further, it is an instructive initial conclusion borne out of an application of the bifurcation/singularity theory steps described in Section 2, that [22]:
@@
The versal unfolding for the OS problem thus requires the inclusion of 3rd-order terms for determinacy, and the problem is of codimension two. We note the planar parameter spaces existing here in the OS problem, rather than the parameter lines we met in the Friedmann problem in Fig. 1, and the two parameters, namely, deviation from spherical symmetry and rotation, make up the versal unfolding and bifurcation diagrams for the deformations. The latter are shown in the two figures 2, 3 below corresponding to positive and negative modular coefficients respectively (cf. [22], Fig. 13, 14).
In these bifurcation diagrams, we imagine a point moving in parameter space on a small circle around the origin and consider the corresponding motion of a phase point moving on some orbit in one of the phase spaces shown. As the parameter point moves around in parameter space, the phase portraits smoothly transfigure to one another, and the phase point drags away from one phase diagram to the next gradually passing through the various metamorphoses of the different topological states (i.e., phase portraits in this case) as shown.
Figure 2: The bifurcation diagram for the Oppenheimer-Snyder-system, positive moduli case, depicting the centre-manifold(s)-reduced dynamics. The bifurcations are as follows: there is a pair of pitchfork bifurcations taking the system along the fragments from 1st to 2nd quadrant, being supercritical in direction right to left ($\mu_{2}>0$); and from 4th to 3rd quadrant, being subcritical in direction right to left ($\mu_{2}<0$). Also there is a supercritical Hopf bifurcation dominated by the $y=\dot{x}$ variable taking the system in the $\mu_{1}<0$ half-space from bottom to top, the bifurcating orbit being stable on the horizontal axis. We note that since the system bifurcates, the number of equilibria changes (as shown in this and the other diagrams), thus leading to new phenomena not having a classical analogue. The escaping orbits signifying collapsing states although appearing everywhere they are never actually realized, as the system transfigures to the ‘next’ possible state upon parameter variation.
Figure 3: The bifurcation diagram for the Oppenheimer-Snyder-system, negative moduli case, on the centre manifold. This is a somewhat more complicated case that includes also non-local bifurcations described by saddle connections (not shown here). The local bifurcations are described by a) a pair of pitchfork bifurcations from the 1st to the 2nd quadrant, supercritical in direction right to left ($\mu_{2}>0$), and from the 4th to the 3rd quadrant, subcritical in direction right to left ($\mu_{2}<0$); b) a subcritical Hopf bifurcation on the $\mu_{1}=\mu_{2}$ line, taking the system in the $\mu_{1}>0$ half-space, the bifurcating orbit being stable on the horizontal axis as before. The solutions are characterized by closed orbits and hence stability, being in a sense closer to describing ‘defocusing’ rather than the focusing states of the positive moduli case in Fig. 2.
Three basic features appearing in these diagrams are:
@@
    the formation of stable closed orbits for various values of the parameters in the left part of the diagram in Fig. 2, and generally in Fig. 3.
@@
    upon parameter variation, the singularities that would have possibly resulted from trapped surface formation and the OS solution are stably bypassed in the present situation. The system instead smoothly transfigures its phase portraits as shown, and any phase point is able to follow other orbits thus easily avoiding staying too long on any runaway orbit.
This is a very rich situation with novel features unlike anything in the original OS analysis. We note that there are additional global bifurcations not shown in these figures.
This constitutes a clear prediction of new physics associated with the versal unfolding in this problem. Under parameter variation, a trapped surface may initially form but such a configuration is always unstable and transfigures into something else precisely defined by new states depending on the position of the parameter point in the stratified space (we recall a stratum as a region in parameter space between two consequent bifurcation sets, e.g., between say a bifurcation axis and a bifurcation curve, etc).
While the original Oppenheimer-Snyder study was restricted to a spherically symmetric dust cloud with spatially constant matter density, it has been subsequently generalized to non-constant matter, non-homogeneous dust clouds, scalar fields, etc, with a partial motivation to tackle the cosmic censorship conjecture (see, [33], Section IV.12, and chap. XIII, for a review of this work with many references to the original sources). The results deployed in this Section may have some relevance to the cosmic censorship problem, and we give here a few brief and very speculative thoughts on this connection.
The use of the word ‘generic’ in this context in the relevant literature in relativity usually refers only to solutions (i.e., generic ‘spacetimes’) rather than to the vector field itself (in particular, one considers for instance ‘generic’ initial data sets specified by a 3-metric and extrinsic curvature). Using vector fields rather than simply their solutions for this purpose is, however, what is involved in the standard definition of a generic property in dynamical systems.
Therefore the problem becomes one to see what properties can be shared by a ‘residual subset in the $C^{k}$-topology’ of a set of nearby vector fields, while one usually considers in this context only the ‘zero-parameter’ solutions (that is those sitting at the stratum ‘origin’ of the parameter space in the bifurcation diagram). This is of course directly related to the idea of structural stability as discussed in this work: the notion of ‘stability of solutions under perturbations’ is relevant to structural stability and genericity only for hyperbolic fixed points and closed orbits (and transversal intersections of stable and unstable manifolds in this case), but not generally for gravitationally unstable or degenerate systems possibly containing non-hyperbolic equilibria etc. Armed with the results described in Figures 2, 3 one thus expects that in a versal deformation of the problem of cosmic censorship in this context, the possible instabilities and appearance of non-generic naked singularities will probably be replaced by stable configurations without singularities, transfiguring to one another upon parameter variations.
@@ 3.4 Spacetime metamorphoses
In this problem we discuss evolutionary aspects of the analysis that leads to the singularities in gravitational collapse to a black hole and in a cosmological setting, cf. [22], Sections 5, 6. In terms of the Komar-Landau-Raychaudhuri and the Newman-Penrose-Sachs equations, this analysis constitutes another manifestation of the nature of implications that are brought by bifurcation theory.
@@
Firstly, it allows for a complete analysis of nonlinear feedback and coupling effects between the expansion, shear, and vorticity of a geodesic congruence in spacetime. These effects can be described by taking into full account all degeneracies of this problem. There are two cases, the convergence-shear and the convergence-vorticity systems (subsystems of the Sachs optical equations), and we shall discuss presently a few of the main ideas and results of the first case (this is in fact the only case considered in the classical references).
Secondly, one can now revisit questions concerning the possible existence of a focussing state. In the classical references this question is of course decided based on the analysis of the Raychaudhuri inequality $\rho^{\prime}\geq\rho^{2}$ for the convergence of the spacetime congruence leading to the prediction of spacetime singularities. In the present approach to this problem, we may apply the methodology deployed earlier in Sec. 2 to the convergence-shear and the convergence-vorticity dynamical equations, since these evolutionary models share all those characteristics necessary for a versal analysis. This requires a set of elaborate techniques which point to the formulation of codimension-two problems associated with the original degenerate dynamics. The final bifurcation diagrams are shown in Figs. 4, 5 respectively. We can only give a few hints here about the global dynamical features revealed by these diagrams. Here the parameters $\mu_{1}$ and $\mu_{2}$ are related to the full projections of the Ricci curvature along a timelike vector and the Weyl curvature along a null tetrad (and rotation parameter respectively for the vorticity case), and the phase space variables are the convergence and the shear (resp. vorticity) of the geodesic congruence.
[…]
Figure 5: The complete bifurcation diagram for the convergence-vorticity system. Here we have again the occurrence of the saddle-node and the pitchfork bifurcations similar to those met in Fig. 4, but now with the important difference that a node instead of a saddle is involved. Additionally, in this case, mainly due to the combined effect of the convergence and vorticity, there is a third kind of bifurcation, namely, a Hopf bifurcation dominated by the vorticity $\omega$ taking the system along the fragments $\chi\to E\rho\to\alpha\to\nu\to\eta,$ or, in the opposite direction. In particular, the cycle created in the degenerate Hopf bifurcation on the $E\rho$-axis stabilizes on the $\alpha$-stratum and makes the Hopf bifurcation non-degenerate in the attractive gravity regions (with the energy condition holding), leading to self-sustained oscillations of the $(\rho,\omega)$ solutions. Upon further parameter variation the cycle disappears to become an unstable focus in the $\eta$ stratum.
@@
As the parameter point moves on a small circle around the origin in the plane, the phase portraits smoothly transfigure to one another, and a phase point rides around from one phase portrait to the next. For instance, the transition from $\chi$ to $\varepsilon_{-}$ to $\pi$ to $\tau_{-}$ to $\beta$ strata (or for the opposite direction) in both figures 4, 5 is accompanied by metamorphoses of non-equivalent phase diagrams as shown, containing 0 to 1 to 2 to 3 equilibria respectively (or in reverse).
@@
In the emerging physics of the versal unfolding, standard conclusions based on the Raychaudhuri inequality are replaced by several novel physical effects of which we only note here two: the possibility of spacetime metamorphoses, and the ‘ghost effect’ associated with a square root scaling law. Both these effects point to a forever-existing universe transfiguring from state to state as shown in the bifurcations, and therefore make the existence of an all-encompassing singular state in general relativity a perhaps less plausible conclusion in the sense that it is not a notable feature of the ensuing bifurcations. Had the bifurcation diagram in Fig. 4 contained only the $\chi$-stratum (i.e., only the right half-space), there would be one phase portrait and nothing else, and this clearly points to the results associated with the primary case of the use of the Raychaudhuri inequality. In that case, there would be of course no possibility for a metamorphosis, because the focussing state leading to a singularity would be an inevitable feature of the dynamics (corresponding to the diverging orbits in the $\chi$-phase portrait, the only kind of orbit there).
@@
However, here the situation is rather different: the system exhibits two kinds of bifurcations, namely, a pair of saddle-nodes dominated by the convergence $\rho$, and a pair of pitchfork bifurcations dominated by the shear $\sigma$ (recall that while the saddle-node bifurcation is the basic mechanism by which equilibria are created and destroyed, a pitchfork comes in two types, the supercritical by which an unstable equilibrium slowly decays into two new stable ones, and the subcritical where the system destabilizes). A saddle-node bifurcation takes the system from the region of attractive gravity $\mu_{1}>0$ where all the energy conditions hold, to the repulsive gravity $\mu_{1}<0$, and back, while a pitchfork occurs in the repulsive gravity region of the parameter space (supercritical for $o\to\tau_{+}\to\beta$, subcritical for $\pi\to\tau_{-}\to\beta$ and back). Hence, the evolution of the system will be mainly characterized by the combined effects of these bifurcations. Another notable feature of the bifurcation dynamics, the ghost or ‘bottleneck’ effect, is related to the saddle-node evolution from and to the $\chi$-stratum, the ‘singularity-forming region’. It implies that the system delays its evolution upon entering the $\chi$-stratum with the delay time scaling as $\mu_{1}^{-1/2}$, a parameter-dependence of the time to the ‘singularity’ as a square-root scaling law.